OpenStudy (anonymous):

A line passes through (-1,7) and (2,10). Which is the equation of the line? A.-3x + y = 16 B -3x+y=4 C -x+y=12 D -x+y=8

OpenStudy (anonymous):

well given that this problem is a multiple choice problem there's a brute force method and the actual solving for the equation method, which do you prefer?

OpenStudy (anonymous):

Brute force? lol

OpenStudy (anonymous):

ok well you can just simply plug in those points into the equations to see if they work, after all if they lie within the line they should work in the equation.

OpenStudy (anonymous):

D. -x+y=8?

OpenStudy (anonymous):

yup! that works doesn't it? Would you like to know the other method or are you good?

OpenStudy (anonymous):

The other method would help because I have a few more of these questions! :)

OpenStudy (anonymous):

haha yes it would. Well you probably know the equation of a line as y =mx+b right?

OpenStudy (anonymous):

Mhm. I actually had to double-check because I was graphing each equation to see. My graphing matched up with letter D, but your tip was a lot quicker and easier.

OpenStudy (anonymous):

ahhh I see well there's another form of the line equation given like this: \[y-y _{1} = m(x-x _{1})\] have you seen this equation?

OpenStudy (anonymous):

Mhm, I have a few times.

OpenStudy (anonymous):

well it's called the point-slope form of an equation. And it can be used to find the equation of a line given you know a point and the slope of the line. Do you know how to solve for the slope of a line?